Given a macaroni economic model:

C=200+2/3yd

G=500

I=300

X=100

M=20+1/10y

T=15+1/20y

Required: determine equilibrium income, consumption exp, total tax, the multiplier


1
Expert's answer
2021-03-01T11:38:46-0500

equilibrium income

"Y=C+I+G-M=200+2\/3(Y-T)+I+G-M"

"Y=200+2\/3(Y-(15+1\/20Y))+300+500-(20+1\/10Y)"

"Y=200+2\/3(19Y\/20-15)+780-1Y\/10=38Y\/60+1Y\/10=970"

"7Y\/15=970"

"Y=2078.57"

consumption expenditure

"C=200+2/3Y_d=200+2/3(2078.57-15-1/20\times 2078.57)=200+1959.64=2159.64"

total tax

"T=15+1/20Y=15+1/20\times 2078.57=118.93"

"multiplier=1\/(1-MPC)" ,"MPC=C\/Y=2159.64\/2078.57=1.039"

"Multiplier=1\/(1-1.039)=-1\/0.039=-25.639"